OpenStudy (anonymous):

Let P(x) be a polynomial with non-negative coefficients. Prove that if the inequality P(1/x)P(x)≥1 holds for x=1, then this inequality holds for all positive x. Sorry for reposting. I think the first one got lost in the afternoon torrent of questions. Happy valentines!

OpenStudy (jamesj):

If this is true for x = 1, then what means is \[ P(1)^2 \geq 1 \]

OpenStudy (anonymous):

I think I'm missing something obvious, but how did you come to that conclusion?

OpenStudy (anonymous):

(Incoming facepalm moment, I'm sure...)

OpenStudy (jamesj):

Now consider the function f(x) = P(1/x)P(x) One way to approach this problem would be to show that for x > 1 , f'(x) > 0 and for 0 < x < 1, f'(x) < 0. In which case, 1 is a local and in fact global minimum of f(x) and therefore \[ f(x) \geq 1 \ \ \] for all real numbers.

OpenStudy (anonymous):

Aha! Alright, I see.

OpenStudy (jamesj):

See if that works for you. I haven't worked this through yet, but I imagine it's a good strategy to start with.

OpenStudy (mr.math):

Oh, I'm totally off. For some reason I solved for P(x)+P(1/x) instead of P(x)P(1/x). I will try again :-)

OpenStudy (jamesj):

So we are given that P(x) has all positive coefficients, \[ P(x) = \sum_{i=0}^n a_ix^i \] where \( a_i \geq 0 \). Now it's clear that for \( P'(x) > 0 \) for \(x > 1 \) and that \( P'(1/x) < 0 \) for \( x > 1 \) for similar reasons. So far so good?

OpenStudy (anonymous):

Sorry, JamesJ, I'm not seeing it. :( Also, I get the feeling that MM's "wrong" proof was still the right way to go about it.

OpenStudy (anonymous):

Lol, near simultaneous posts.

OpenStudy (anonymous):

How do we know that a is greater than or equal to 0?

OpenStudy (jamesj):

"Let P(x) be a polynomial with non-negative coefficients"

OpenStudy (anonymous):

Woops. My bad.

OpenStudy (anonymous):

Sorry, this is what happens when I try to multitask. :| Alright, so far so good.

OpenStudy (jamesj):

Now \[ f'(x) = P(1/x)P'(x) - \frac{1}{x^2}P'(1/x)P(x) \] Show now that for \( x > 1 \), \( f'(x) > 0 \)

OpenStudy (jamesj):

And in the meantime, if anyone has a more elegant proof, the floor is yours

OpenStudy (mr.math):

@James: I think your proof doesn't include the interval (0,1), does it?

OpenStudy (jamesj):

No, it doesn't. But show this part and then you can get the other by (anti)symmetry.

OpenStudy (mr.math):

Ah I see.

OpenStudy (anonymous):

I get it so far. But what do you mean by "antisymmetry"?

OpenStudy (mr.math):

I think he meant something like this: for every number \(a>1\), there exist a number \(b<1\) such that \(P(a)P(\frac{1}{a})=P(b)P(\frac{1}{b})\), namely \(b=\frac{1}{a}\).

OpenStudy (anonymous):

Huh. While I mull over the implications of that, I'd like to point out the "elegant" proof for this is as attached.

OpenStudy (anonymous):

I didn't come up with it, naturally. XD

OpenStudy (jamesj):

Right.

OpenStudy (anonymous):

Solving it without Cauchy would be cool too, though.

OpenStudy (mr.math):

I don't really know this inequality. So for me I would have to prove this inequality first.

OpenStudy (jamesj):

I'm sure you do it, just that it's disguised here. The usual form is this: consider any two vectors v and w in \( \mathbb{R}^n \). Then \[ u \cdot v \leq ||u|| \ ||v|| \]

OpenStudy (anonymous):

Man, I'm confused. What course am I supposed to learn this in?

OpenStudy (anonymous):

Oh, wait, nevermind. I'm not confused. XD I type faster than I think.

OpenStudy (mr.math):

lol.

OpenStudy (jamesj):

You're meta-confused: confused about your state of confusion.

OpenStudy (anonymous):

Haha.

OpenStudy (anonymous):

Oh, by the way, the non-Cauchy solution as given. Attached. Naturally, again, I was not involved in doing this. :P Although this one makes a lot more sense to me.

OpenStudy (anonymous):

Now that I think about it, though, that middle looks an awful lot like Cauchy...

OpenStudy (mr.math):

This last proof is what I've been trying to do. I like it more than the others.

OpenStudy (jamesj):

For what it's worth, to finish off my first strategy: If P(1)^2 = 1, then P(1) = 1 (as all the coefficients are positive) and P(x) > 0 for all x > 1; similarly P(1/x) < 1 for all x > 1. Therefore f'(x) = P'(x)P(1/x) - (1/x^2)P(x)P'(1/x) > P'(x).1 - 1.1.P'(1/x) > 0 Now, for x > 1, 0 < 1/x < 1, and \[ (f(1/x))' = -(1/x^2)f'(1/x) = -(1/x^2)f'(x) \] because f(x) = f(1/x) Now that last expression is less than zero because f'(x) > 0 for x > 1. Hence f has the properties we want and hence the inequality is proved. === The C-S proof is still better. Much better. And no, there second proof doesn't use it, but another calculus result instead: that for any positive integer n \[ x^n + \frac{1}{x^n} \geq 2 \]

OpenStudy (jamesj):

*their second proof

OpenStudy (anonymous):

I found your proof the easiest to understand. :P Thanks for the help. I was hoping to derive another method of proving it, other than the two I provided.